OpenStudy (anonymous):

Find a function that models the surface area S of a cube in terms of its volume V. I know the surface area = 6x^2 volume = x^3 S(V)= ???

jimthompson5910 (jim_thompson5910):

Volume: V = x^3 solve for x to get x = ??

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} &1ft & = & 12in\\ &1ft^2 & = & 12^2in^2\\ &1ft^3 & = & 12^3in^3\\ \textit{thus we can say}\\ &area & = & x\cdot x\\ &volume& = & x\cdot x \cdot x\\ \hline\\ S(v) = \cfrac{\sqrt{6}\cdot \sqrt{6} \cdot \sqrt{6}x^3}{\sqrt{6} x} \implies 6x^2 \end{array}\)

OpenStudy (anonymous):

Hmm, still says it's wrong. I just don't understand q.q

jimthompson5910 (jim_thompson5910):

what did you type in exactly?

OpenStudy (anonymous):

6x^2

jimthompson5910 (jim_thompson5910):

that's the surface area, but it's in terms of the side length and NOT the volume

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."